OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (950, 1,250) and has a slope of –10?

OpenStudy (texaschic101):

y = mx + b slope(m) = -10 (950,1250)...x = 950 and y = 1250 now we sub 1250 = -10(950) + b 1250 = -9500 + b 1250 + 9500 = b 10750 = b your equation is : y = -10x + 10,750

OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)?